# Distributed Online Economic Dispatch With Time-Varying Coupled Inequality Constraints

### v1

@@
Keywords: Distributed decision making, Economic dispatch, Distributed online optimization, Time-varying coupled inequality constraints, Optimal decisions.
@@ 1 Introduction
where $\Omega_{i}\subset\mathbb{R}$ denotes the power output range of generator $i$, $P_{t}$ is the real-time electricity price, $a_{i,t},b_{i,t},c_{i,t}$ are the generation cost coefficients, and $D_{t}$ is the system load demand at time $t$. The term $P_{t}x_{i,t}$ in the objective function represents the revenue or market settlement at the real‑time price, which is subtracted from the generation cost to compute the net operating cost. The coupled inequality constraint $D_{t}-\sum_{i=1}^{N}x_{i,t}\leq 0$ ensures that the total generation reaches or exceeds the system load at each time step. Since the generation cost in (1.1) is increasing and strongly convex in $x_{i,t}$, any surplus generation increases the total cost. Hence, the optimal solution satisfies the coupled constraint at equality, i.e., $D_{t}-\sum_{i=1}^{N}x_{i,t}=0$. Note that, at the decision-making stage, both $P_{t}$ and $D_{t}$ are time-varying and unobserved, and thus are replaced by their forecasts. Moreover, each unit can typically observe only its own cost and constraint information, without direct knowledge of the global demand or the private parameters of other units, making the problem highly challenging. Existing methods for problem (1.1) are largely heuristic and lack rigorous theoretical analysis (Guo2007; Sunwen2022); furthermore, some studies consider illustrative examples that do not arise from practical engineering scenarios, limiting their practical applicability (xiesiyu2022; Lijueyou2022).
@@ 1.1 Related works
In fact, (1.1) can be formulated as a distributed online optimization problem. In recent years, driven by rapid advances in communication and computation technologies, together with the widespread deployment of low‑cost devices, distributed optimization has attracted substantial attention across a broad range of research areas. Many practical application scenarios, such as decision making, statistical learning, sensor networks, resource allocation, and formation control (Abouarghoub2018; ChenT2017; MaW2018; Bedi2018), can be naturally formulated as distributed optimization tasks over multi-agent networks. Unlike traditional centralized optimization methods, distributed optimization is performed on a network of agents, each of which has access only to its private information, while no central coordinator possesses complete system knowledge. Consequently, no single agent can solve the global optimization problem on its own; instead, the global objective can be achieved only through information exchange and coordinated decision-making among agents.
@@ 1.3 Brief discussion
where $x\in\mathbb{R}^{p}$ is the common decision vector, $H_{i,t}\in\mathbb{R}^{d_{i}\times p}$, $h_{i,t}\in\mathbb{R}^{d_{i}}$, and $b_{i,t}\in\mathbb{R}$ with $d_{i}\in\mathbb{N}_{+}$.
@@ 2.2 Problem formulation
where $f_{i,t}:\Omega_{i}\rightarrow\mathbb{R}$ is the local objective function of node $i$ at time $t$, $g_{i,t}:\Omega_{i}\rightarrow\mathbb{R}^{n}$ is the local constraint function of node $i$ at time $t$, $x_{i}\in\Omega_{i}\in\mathbb{R}^{d_{i}}$ is the decision variable of node $i$, $x=(x_{1}^{\top},\dots,x_{N}^{\top})^{\top}\in\mathbb{R}^{d}$, $\sum_{i=1}^{N}d_{i}=d$, and $T$ is the time horizon. At each time $t$, node $i$ only knows $f_{i,t}$, $g_{i,t}$ and $\Omega_{i}$. Let $\Omega=\Omega_{1}\times\dots\times\Omega_{N}$. The feasible set of problem (2.1) at time $t$ is $\bar{\Omega}_{t}=\{x\in\Omega|g_{t}(x)\leq\mathbf{0}_{n}\}$.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No. 62261136550 and the NYU-ECNU Institute of Mathematical Sciences, NYU Shanghai, Shanghai, China.

### v2

@@
Keywords: Optimal decisions, Economic dispatch, Distributed online optimization, Time-varying coupled inequality constraints, Power systems.
@@ 1 Introduction
where $\Omega_{i}\subset\mathbb{R}$ denotes the power output range of generator $i$, $P_{t}$ is the real-time electricity price, $a_{i,t},b_{i,t},c_{i,t}$ are the generation cost coefficients, and $D_{t}$ is the system load demand at time $t$. The term $P_{t}x_{i,t}$ in the objective function represents the revenue or market settlement at the real‑time price, which is subtracted from the generation cost to compute the net operating cost. The coupled inequality constraint $D_{t}-\sum_{i=1}^{N}x_{i,t}\leq 0$ ensures that the total generation reaches or exceeds the system load at each time step. Since the generation cost in (1.1) is increasing and strongly convex in $x_{i,t}$, any surplus generation increases the total cost. Hence, the optimal solution satisfies the coupled constraint at equality, i.e., $D_{t}-\sum_{i=1}^{N}x_{i,t}=0$. Note that, at the decision-making stage, both $P_{t}$ and $D_{t}$ are time-varying and unobserved, and thus are replaced by their forecasts. Moreover, each unit can typically observe only its own cost and constraint information, without direct knowledge of the global demand or the private parameters of other units, making the problem highly challenging. Existing methods for problem (1.1) are largely heuristic and lack rigorous theoretical analysis (Guo2007; cortes2018genetic; Sunwen2022); furthermore, some studies consider illustrative examples that do not arise from practical engineering scenarios, limiting their practical applicability (xiesiyu2022; Lijueyou2022).
@@ 1.1 Related works
In fact, (1.1) can be formulated as a distributed online optimization problem. In recent years, driven by rapid advances in communication and computation technologies, together with the widespread deployment of low‑cost devices, distributed optimization has attracted substantial attention across a broad range of research areas. Many practical application scenarios, such as decision making, statistical learning, sensor networks, resource allocation, and formation control (luo2023optimal; Abouarghoub2018; ChenT2017; MaW2018; Bedi2018), can be naturally formulated as distributed optimization tasks over multi-agent networks. Unlike traditional centralized optimization methods, distributed optimization is performed on a network of agents, each of which has access only to its private information, while no central coordinator possesses complete system knowledge. Consequently, no single agent can solve the global optimization problem on its own; instead, the global objective can be achieved only through information exchange and coordinated decision-making among agents.
@@ 1.3 Brief discussion
where $x\in\mathbb{R}^{p}$ is the common decision vector, $H_{i,t}\in\mathbb{R}^{d_{i}\times p}$, $h_{i,t}\in\mathbb{R}^{d_{i}}$, $b_{i,t}\in\mathbb{R}$, and $d_{i}\in\mathbb{N}_{+}$ represents the output/response dimension of node $i$.
@@ 2.2 Problem formulation
where $f_{i,t}:\Omega_{i}\rightarrow\mathbb{R}$ is the local objective function of node $i$ at time $t$, $g_{i,t}:\Omega_{i}\rightarrow\mathbb{R}^{n}$ is the local constraint function of node $i$ at time $t$, $x_{i}\in\Omega_{i}\in\mathbb{R}^{d_{i}}$ is the decision variable of node $i$, $d_{i}$ denotes the dimension of the decision variable of node $i$, $x=(x_{1}^{\top},\dots,x_{N}^{\top})^{\top}\in\mathbb{R}^{d}$, $\sum_{i=1}^{N}d_{i}=d$, and $T$ is the time horizon. At each time $t$, node $i$ only knows $f_{i,t}$, $g_{i,t}$ and $\Omega_{i}$. Let $\Omega=\Omega_{1}\times\dots\times\Omega_{N}$. The feasible set of problem (2.1) at time $t$ is $\bar{\Omega}_{t}=\{x\in\Omega|g_{t}(x)\leq\mathbf{0}_{n}\}$.
@@ Acknowledgments
Tao Li was supported by the National Natural Science Foundation of China under Grant No. 62261136550 and the NYU-ECNU Institute of Mathematical Sciences, NYU Shanghai, Shanghai, China. Xiaoqian Wang was supported by the Presidential Foundation of the Academy of Mathematics and Systems Science, Chinese Academy of Sciences, China (No. E555930101).